OpenStudy (anonymous):

The ratio of the heights of two similar rectangular prisms is 3:4. What is the ratio of their lateral areas? Answer 3:7 1:0.75 9:16 27:64

OpenStudy (anonymous):

lateral surface = l*h: the prisms are similar implies that h1/h2=l1/l2=3/4 and what we find that area1/area2=l1*h1/l2*h2=3*3/4*4 =9/16

OpenStudy (anonymous):

thank-you!!